OpenStudy (rainbow_rocks03):

If ∠A and ∠B are supplementary angles and if m∠A is 68°, what is m∠B? A. 112° B. 22° C. 136° D. 68°

jimthompson5910 (jim_thompson5910):

A & B are supplementary, so A+B = 180

jimthompson5910 (jim_thompson5910):

A+B = 180 68 + B = 180 B = ???

OpenStudy (anonymous):

A, 112

OpenStudy (rainbow_rocks03):

No direct answers !

OpenStudy (anonymous):

well sorry

OpenStudy (rainbow_rocks03):

thx for the help @jim_thompson5910

jimthompson5910 (jim_thompson5910):

np